Prove or disprove:

a.) If 𝑃(𝐴) = 𝑃(𝐡) = 𝑝 then 𝑃(𝐴𝐡) ≀ 𝑝^2 .

b.) If 𝑃(𝐴) = 0 then 𝑃(𝐴𝐡) = 0.

c.) If 𝑃(𝐴̅) = π‘Ž and 𝑃(𝐡̅) = 𝑏 then 𝑃(𝐴𝐡) β‰₯ 1 βˆ’ π‘Ž βˆ’ 𝑏.


1
Expert's answer
2021-03-19T15:53:11-0400

Prove or disprove:

a.) If 𝑃(𝐴) = 𝑃(𝐡) = 𝑝 then 𝑃(𝐴𝐡) ≀ 𝑝^2 .

Put B=A, then 𝑃(𝐴) = 𝑃(𝐡) = 𝑝, but 𝑃(𝐴𝐡)=𝑃(𝐴) =𝑝> 𝑝^2 (if p is not equal to 0 or 1). The assertion is disproved.


b.) If 𝑃(𝐴) = 0 then 𝑃(𝐴𝐡) = 0.

This is true, as "AB\subset A" and 𝑃(𝐴𝐡)≀𝑃(𝐴)=0. Therefore 𝑃(𝐴𝐡)=0.


c.) If 𝑃(𝐴̅) = π‘Ž and 𝑃(𝐡̅) = 𝑏 then 𝑃(𝐴𝐡) β‰₯ 1 βˆ’ π‘Ž βˆ’ 𝑏.

False. If we take a=b=0, we would get 𝑃(𝐴𝐡) β‰₯ 1 which contradicts to the inequality 𝑃(𝐴𝐡)≀𝑃(𝐴)=0.